# ElastoGen: 4D Generative Elastodynamics

### v2

@@ Abstract
1 Introduction
Recent advancements in generative models have enhanced the ability to produce high-quality digital contents across diverse media formats (e.g. images, videos, 3D models, 4D data). In particular, the generation of 4D data, including both spatial and temporal dimensions, has seen notable progress (Singer et al., 2023; Shen et al., 2023; Xu et al., 2024; Ling et al., 2023; Bahmani et al., 2024a; Yin et al., 2023; Bahmani et al., 2024b).
@@ 1 Introduction
We explore a new way to establish physics-in-the-loop generative models. Our argument is that learning from knowledge instead of from raw data is more effective for generative models. Physical laws and principles are often in the form of partial differential equations (PDEs) and numerically solved with discretized differential operators. We note that those operators hold a similar structure as a convolution kernel on the problem domain, where the values of those convolution kernels depend on the specific problem setting. Inspired by those observations, we propose ElastoGen, a knowledge-driven neural model that generates physically accurate and coherent 4D elastodynamics. ElasoGen can be easily coupled and integrated with upstream and downstream neural modules to enable end-to-end 4D generation. The core idea of ElastoGen is converting the global differential operator, corresponding to the nonlinear force equilibrium, to iterative local convolution-like procedures. Such knowledge-level priors allow us to design dedicated network modules for ElastoGen, where each network module has a well-defined purpose of relaxing locally concentrated strain rather than being treated as a piece of a “black box”. Compared with other data-learning-based generative models, ElasoGen is lightweight – in terms of both training requirements and the network scales. Furthermore, due to its consistency with physics procedure, ElastoGen generates physically accurate dynamics for a wide range of hyperelastic materials. Specifically, we summarize some features of ElastoGen as follows:
@@
NeuralMTL with diffusion parameterization   ElastoGen features a so-called NeuralMTL module to encode the underlying constitutive relations for real-world hyperelastic materials such as Neo-Hookean and or Saint Venant-Kirchhoff (StVK). We leverage a lightweight conditional diffusion model to predict its network parameters to isolate our training efforts.
@@
Nested RNN with low-frequency encoding   ElastoGen constitutes a two-level RNN architecture. We augment ElastoGen with a low-frequency encoder, which extracts low-frequency dynamic signals so that the local relaxation only takes care of the remaining high-frequency strains. This design makes ElastoGen more efficient for stiff instances.
2 Related Work
Generative models The primary objective of generative models is to produce new, high-quality samples from vast datasets. These models are designed to learn and understand the distribution of data, thereby generating samples that meet specific criteria. Techniques such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Variational Autoencoders (VAEs) (Kingma & Welling, 2014), and flow-based methods (Dinh et al., 2015; 2017) have all demonstrated significant success. However, each method has its limitations. For instance, GANs can generate high-quality images but are notoriously difficult to train and optimize (Arjovsky et al., 2017; Gulrajani et al., 2017; Mescheder, 2018). VAEs (Vahdat & Kautz, 2020; Child, 2021) and flow-based methods (Kingma & Dhariwal, 2018) offer efficient training processes but generally fall short in sample quality compared to GANs. Recently, diffusion models have emerged as another powerful technique, achieving state-of-the-art results in generating high-fidelity images (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022), setting the stage for further explorations in more complex applications.
@@ 2 Related Work
4D generation based on diffusion models As research on diffusion models advances, these methods could potentially be applied to the generation of 3D content (Jain et al., 2022; Lin et al., 2023; Metzer et al., 2023; Poole et al., 2022; Wang et al., 2024b; Liu et al., 2023; 2024), video content (Blattmann et al., 2023; Harvey et al., 2022; Ho et al., 2022b; a; Karras et al., 2023; Ni et al., 2023), and more complex forms such as 3D videos or what might be termed 4D scenes (Singer et al., 2023; Shen et al., 2023; Xu et al., 2024; Ling et al., 2023; Bahmani et al., 2024a; Yin et al., 2023; Bahmani et al., 2024b). These advanced applications demonstrate the versatility and expanding potential of diffusion models across diverse media formats. However, existing video generation techniques struggle to ensure temporal consistency and require substantial training data, underscoring the challenges of capturing and replicating the dynamic and interconnected behaviors present in real-world scenarios within a generative model framework.
@@
Neural physical synamics Physical dynamics traditionally relies on numerical solutions such as the finite element method (FEM) (Zienkiewicz & Morice, 1971; Zienkiewicz et al., 2005; Huebner et al., 2001; Reddy, 1993), finite difference method (Zhu et al., 2010; Godunov & Bohachevsky, 1959), or mass-spring systems (Liu et al., 2013). Each approach offers distinct advantages and limitations. For example, Position-Based Dynamics (PBD) (Müller et al., 2007) and Projective Dynamics (PD) (Bouaziz et al., 2014; Liu et al., 2013) offer simplified implementation and faster convergence but can struggle with complex material behaviors and do not always guarantee consistent convergence rates. Recently, neural physics solvers, which integrate neural networks with traditional solvers, aim to accelerate and simplify the computation process. The pioneering works (Chang et al., 2017; Battaglia et al., 2016) directly utilized neural networks to predict dynamics, achieving promising results in simple particle systems. Subsequent studies (Sanchez-Gonzalez et al., 2018; Kipf et al., 2018; Ajay et al., 2018; Li et al., 2019c; a; b) adopted network architectures to the specific features of the systems, thereby enhancing performance. The advent of Physics Informed Neural Networks (PINNs) (Raissi et al., 2019; Pakravan et al., 2021) marks a leap forward. These networks incorporate extensive physical information to constrain and guide the learning process, ensuring that predictions adhere more closely to physical laws and has succeeded in domains such as cloths (Geng et al., 2020) and fluids (Um et al., 2020; Gibou et al., 2019; Chu et al., 2022). Some work (Yang et al., 2020) shifts away from end-to-end structures and use neural networks to optimize part of the simulation. Another line of research generates dynamics through physics-based simulators, where network learns static information while physical laws govern the generation of dynamics (Li et al., 2023; Feng et al., 2023; Xie et al., 2023; Feng et al., 2024; Jiang et al., 2024), giving physical meanings to Neural Radiance Fields (NeRF) (Mildenhall et al., 2020; Kerbl et al., 2023a). These methods demonstrate the benefits of embedding human knowledge into networks to reduce the learning burden.
3 Background
@@
To make the paper more self-contained, we start with a brief review of some preliminaries of a dynamic elastic model.
3.1 Variational optimization of elastodynamics
@@
The dynamic equilibrium of a 3D model can be characterized by $\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L}{\partial\dot{\mathbf{q}}}\right)-\frac{\partial L}{\partial\mathbf{q}}=\mathbf{f}_{q}$, where $L=T-U$ is system Lagrangian i.e., the difference between the kinematic energy ($T$) and the potential energy ($U$). $\mathbf{q}$ and $\dot{\mathbf{q}}$ are generalized coordinate and velocity. $\mathbf{f}_{q}$ is the generalized external force. With the implicit Euler time integration scheme: $\mathbf{q}_{n+1}=\mathbf{q}_{n}+h\dot{\mathbf{q}}_{n+1}$, $\dot{\mathbf{q}}_{n+1}=\dot{\mathbf{q}}_{n}+h\ddot{\mathbf{q}}_{n+1}$, it can be reformulated as a nonlinear optimization to be solved at each time step:
$\mathbf{q}_{n+1}=\underset{\mathbf{q}}{\mathrm{argmin}}\left\{\frac{1}{2h^{2}}\|\mathbf{q}-\mathbf{q}_{n}-h\dot{\mathbf{q}}_{n}-h^{2}\mathbf{M}^{-1}\mathbf{f}_{q}\|^{2}_{\mathbf{M}}+U(\mathbf{q})\right\},$ (1)
@@
where the subscript indicates the time step index. $h$ is the time step size, and $\mathbf{M}$ is the mass matrix.
3.2 Diffusion model
@@
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{(1-\overline{\alpha}_{t})}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t))+\sigma_{t}\mathbf{z},\quad\mathbf{z}\sim N(\mathbf{0},\mathbf{I}),$ (3)
@@
4 Methodology
Our overall pipeline is visualized in figure 1. ElastoGen is a lightweight generative deep model producing physically grounded 4D contents given some general descriptions of the object e.g., stiffness or density.Such information could also be learned via observations since ElastoGen is trivially differentiable. ElastoGen rasterizes an input shape and leverages a nested two-level RNN to predict its further trajectory sequentially. Each prediction is subject to an accuracy check to ensure the result is physically accurate. Such network structure adheres to a well-reasoned numerical procedure for solving the variational optimization of equation 1. Therefore, ElastoGen does not have redundant or purpose-less network components that could potentially lead to overfitting. In the following sections, we elaborate on each major module of our pipeline.
@@ 4 Methodology
4.1 Method overview: piece-wise local quadratic approximation
@@
Our elastodynamic generation mimics numerical optimization procedures that minimize the variational energy of equation 1. It is possible to tackle this problem at the global level, i.e., optimizing all the degrees of freedom (DoFs) of the system at once e.g., using Newton’s method. Such a brute-force scheme requires to learn dense inter-correlations among features at all DoFs, which inevitably leads to complex and large-scale network architectures with numerous parameters to be learned.
@@
Alternatively, we opt for a divide-and-conquer way to approach equation 1. We consider the total potential energy $U$ as the summation of multiple energies of quadratic form: $U(\mathbf{q})\approx\sum_{i}E_{i}(\mathbf{q}_{i})$, where $E_{i}(\mathbf{q})=\min_{\mathbf{p}_{i}\in\mathcal{M}_{i}}\frac{\omega_{i}}{2}\left\|\mathcal{G}_{i}[\mathbf{q}_{i}]-\mathbf{p}_{i}\right\|^{2}$. Here, $i$ indicates the $i$-th sub-volume of the object. For instance, one may discretize the object into a tetrahedral mesh, and $E_{i}$ then represents the elastic potential stored at the $i$-th element. $\mathcal{G}_{i}$ denotes a discrete differential operator, which converts positional features $\mathbf{q}_{i}$ to strain-level features. To this end, we build $\mathcal{G}_{i}$ such that $\mathcal{G}_{i}[\mathbf{q}_{i}]=\text{vec}(\mathbf{F}_{i})$, i.e., the vectorized deformation gradient ($\mathbf{F}\in\mathbb{R}^{3\times 3}$) of the sub-volume, which gives the local first-order approximation of the displacement field. The constraint manifold $\mathcal{M}_{i}$ denotes the zero level set of $E_{i}$. In other words, we consider $E_{i}$ as a quadratic energy based on how far local displacement $\mathbf{q}_{i}$ is from its closest energy-free configuration ($\mathbf{p}_{i}$), given the local material stiffness $\omega_{i}$.
@@
Provided the current deformed shape $\mathbf{q}_{i}$, we can find $\arg\min_{\mathbf{p}_{i}}\frac{\omega_{i}}{2}\left\|\mathcal{G}_{i}[\mathbf{q}_{i}]-\mathbf{p}_{i}\right\|^{2}$, which suggests a locally optimal descent direction to reduce $U$. The global displacement can then be obtained by minimizing $\mathbf{q}$ over $E_{i}$ at all the sub-volumes. While this is a global operation that we would like to avoid, it is essentially a Laplacian-like smoothing operator, which can still be processed with repeated local smoothing. This procedure share a similar nature of shape matching method (Müller et al., 2005) and PD (Bouaziz et al., 2014) — it offers a piece-wise SQP way (Boggs & Tolle, 1995) to approximate $U$ locally. ElastoGen functions like a neural version of the aforementioned procedure with a nested RNN structure. It handles local solve, or strain relaxation in the form of a volume convolution so that the overall network structure is compact and lightweight.
@@
Unfortunately, real-world materials are more than a collection of quadratic forms. The appropriate $\mathcal{M}_{i}$ nonlinearly vary under different deformation or material models, aka material nonlinearity. As a result, shape matching or PD can only handle simplified material behavior unless we know how $\mathcal{M}_{i}$ changes along the generation. To this end, we augment ElastoGen with a NeuralMTL module to make sure each local SQP matches actual materials.
4.2 NeuralMTL & neural projection
@@
The goal of NeuralMTL is to correct local quadratic approximations of $U$ so that ElastoGen faithfully generates physically accurate results for any real-world hyperelastic material. Specifically with NerualMTL ($\mathcal{N}$), $E_{i}$ becomes:
@@
We set $\omega_{i}$ as $\omega_{i}=V_{i}{e}/({2(1+\nu)})$ based on real-world material parameters: Young’s modulus $e$, Poisson’s ratio $\nu$ as well as the size of the sub-volume $V_{i}$. $\mathcal{G}_{i}$ extracts the deformation gradient $\text{vec}(\mathbf{F}_{i})$ and feeds it to NeuralMTL, $\mathcal{N}$. As the name suggests, $\mathcal{N}$ predicts a neural strain based on the information of local deformation $\mathbf{F}_{i}$. Given the material model and parameters, $\mathcal{N}$ is used for all $E_{i}$, and we do not put a subscript on $\mathcal{N}$. $\|\cdot\|_{F}$ denotes the Frobenius norm. $\mathcal{N}$ predicts a material-space strain prediction, which is then converted to is world space by $\mathbf{F}_{i}$. $\mathbf{P}_{i}\in\mathbb{R}^{3\times 3}$ is a rotation matrix i.e., $\mathbf{P}_{i}\in\mathcal{SO}(3)$. Intuitively, NeuralMTL warps $\mathbf{F}_{i}$ to a different configuration of $\mathbf{F}_{i}\cdot\mathcal{N}(\mathbf{F}_{i})$ so that the new distance to $\mathbf{P}_{i}$ correctly reflects the local energy landscape of $E_{i}$ as visualized in the left inset.
For isotropic elastic materials, we add a nonlinear SVD (singular value decomposition) activation to the operator $\mathcal{G}_{i}$ such that $\mathbf{F}_{i}=\mathbf{U}_{I}\mathbf{S}_{I}\mathbf{V}_{i}^{\top}$. $\mathbf{S}_{i}$ is a diagonal matrix with singular values arranged in descending order, which correspond to the local principal strains. This activation converts $E_{i}$ to:
$\begin{split}E_{i}(\mathbf{q}_{i})&=\frac{\omega_{i}}{2}\|\mathbf{U}_{i}\mathbf{S}_{i}\mathbf{V}_{i}^{\top}\cdot\mathcal{N}(\mathcal{G}_{i}[\mathbf{q}_{i}])-\mathbf{U}_{i}\mathbf{V}_{i}^{\top}\|^{2}_{F}\\ &=\frac{\omega_{i}}{2}\mathrm{tr}\left(\mathbf{S}_{i}\mathbf{S}_{i}\mathbf{V}_{i}^{\top}\cdot\mathcal{N}(\mathcal{G}_{i}[\mathbf{q}_{i}])\cdot\mathcal{N}^{\top}(\mathcal{G}_{i}[\mathbf{q}_{i}])\mathbf{V}_{i}+\mathbf{I}-2\mathbf{V}_{i}\mathbf{S}_{i}\mathbf{V}_{i}^{\top}\cdot\mathcal{N}(\mathcal{G}_{i}[\mathbf{q}_{i}])\right).\end{split}$ (5)
@@
$\begin{split}E_{i}=\frac{\omega_{i}}{2}\mathrm{tr}\left(\mathbf{Q}_{i}\mathbf{Q}_{i}^{\top}\right)+\frac{3\omega_{i}}{2}-\omega_{i}\ \mathrm{tr}\left(\mathbf{Q}_{i}\right),\;\mathbf{Q}_{i}(\mathbf{S}_{i})=\mathbf{S}_{i}\left(\mathbf{N}_{i}+\mathbf{N}_{i}^{\top}\right).\end{split}$ (7)
Intuitively, this activation escalates the order of the neural strain predicted by $\mathcal{N}$, pushing it to become a nonlinear strain estimation with a prescribed format — just like upgrading an infinitesimal strain to Green’s strain to better measure large rotational deformation. It should be noted that NerualMTL prediction not alter the location of $\mathbf{P}_{i}$. As a result, the neural projection corresponding to our NerualMTL can be easily obtained as $\mathbf{P}_{i}=\mathbf{U}_{i}\mathbf{V}_{i}^{\top}$, i.e., the rotational component from $\mathbf{F}_{i}$. This is an important property of NerualMTL — if we choose to employ the network to learn an adjustment of $\mathbf{P}_{i}$ (which is also technically feasible), the local relaxation that predicts $\mathbf{P}_{i}$ becomes complicated, and the generation is less robust.
Given an input 3D object, ElastoGen rasterizes it into a set of 3D voxels. For a user-specified sub-volume e.g., in our implementation, each sub-volume is a voxel that intersects with the object, $\mathcal{G}_{i}$ operator extracts the local covariance matrix of the displacement field over this sub-volume. Let $\mathbf{A}_{i}=[\mathbf{q}_{1},\mathbf{q}_{2},...\mathbf{q}_{k}]\in\mathbb{R}^{3\times k}$ and $\bar{\mathbf{A}}_{i}=[\bar{\mathbf{q}}_{1},\bar{\mathbf{q}}_{2},...\bar{\mathbf{q}}_{k}]$ be deformed and rest-shape position of vertices of a sub-volume with $k$ vertices ($k=8$ for a cubic volume). $\mathcal{G}_{i}$ has an analytic format of:

### v3

@@ Abstract
Introduction
Recent advancements in generative models have enhanced the ability to produce high-quality digital contents across diverse media formats (e.g. images, videos, 3D models, 4D data). In particular, the generation of 4D data, including both spatial and temporal dimensions, has seen notable progress (singer2023text; shen2023make; xu2024comp4d; ling2023align; bahmani2024tc4d; yin20234dgen; bah20244dfy).
@@ Introduction
We explore a new way to establish physics-in-the-loop generative models. Our argument is that learning from knowledge instead of from raw data is more effective for generative models. Physical laws and principles are often in the form of partial differential equations (PDEs) and numerically solved with discretized differential operators. We note that those operators hold a similar structure as a convolution kernel on the problem domain, where the values of those convolution kernels depend on the specific problem setting. Inspired by those observations, we propose ElastoGen, a knowledge-driven neural model that generates physically accurate and coherent 4D elastodynamics. ElastoGen can be easily coupled and integrated with upstream and downstream neural modules to enable end-to-end 4D generation. The core idea of ElastoGen is converting the global differential operator, corresponding to the nonlinear force equilibrium, to iterative local convolution-like procedures. Such knowledge-level priors allow us to design dedicated network modules for ElastoGen, where each network module has a well-defined purpose of relaxing locally concentrated strain rather than being treated as a piece of a black box. Compared with other data-learning-based generative models, ElastoGen is lightweight – in terms of both training requirements and the network scales. Furthermore, due to its consistency with physics procedure, ElastoGen generates physically accurate dynamics for a wide range of hyperelastic materials. Specifically, we summarize some features of ElastoGen as follows:
@@
NeuralMTL with diffusion parameterization  ElastoGen features a neural material module, NeuralMTL, to encode the underlying constitutive relations for real-world hyperelastic materials such as Neo-Hookean and or Saint Venant-Kirchhoff (StVK). We leverage a lightweight conditional diffusion model to predict its network parameters to isolate our training efforts.
@@
Nested RNN with low-frequency encoding  ElastoGen constitutes a two-level recurrent neural network (RNN) architecture. An encoder extracts low-frequency deformations so that the inner RNN relaxation only takes care of the local high-frequency strains. This design makes ElastoGen more efficient for stiff materials.
Related work
@@
4D generation with diffusion models The primary objective of generative models is to produce new, high-quality samples from vast datasets. These models are designed to learn and understand the distribution of data, thereby generating samples that meet specific criteria. Recently, diffusion models have emerged as a powerful technique, achieving state-of-the-art results in generating high-fidelity images (Sohl-DicksteinW2015diffusion; HoJA2020ddpm; rombach2022high), beating the competitors that are based on generative adversarial networks (GANs) (goodfellow2014gan; zhu2019dm; chan2021pi) or variational autoencoders (VAEs) (kingma2014vae; Child2021vae; razavi2019generating). This sets the stage for further explorations in more complex applications, such as the generation of 3D content (jain2022zero; lin2023magic3d; metzer2023latent; poole2022dreamfusion; wang2024prolificdreamer; liu2023zero; liu2023one2345; feng2024arm), video content (blattmann2023align; harvey2022flexible; ho2022video; ho2022imagen; karras2023dreampose; ni2023conditional), and 3D videos or called 4D dynamic scenes (3D objects with shape change throughout time) (singer2023text; shen2023make; xu2024comp4d; ling2023align; bahmani2024tc4d; yin20234dgen; bah20244dfy). These advanced applications demonstrate the versatility and expanding potential of diffusion models across diverse media formats. However, existing 4D generation techniques struggle to ensure temporal consistency and require substantial training data, underscoring the challenges of capturing and replicating the dynamic and interconnected behaviors present in real-world scenarios within a generative model framework.
@@
Neural physical dynamics Physical dynamics has traditional numerical solutions, such as the finite element method (FEM) (zienkiewicz1971finite; zienkiewicz2005finite) and mass-spring systems (liu2013pd). Each approach offers distinct advantages and limitations. For example, Position-Based Dynamics (PBD) (muller2007pbd) and Projective Dynamics (PD) (Bouaziz2014PD) offer simplified implementation and faster convergence but can struggle with complex material behaviors and do not always guarantee consistent convergence rates. Recently, neural physics solvers, which integrate neural networks with traditional solvers, aim to accelerate and simplify the computation process. The pioneering works (chang2017object; battaglia2016interaction) directly utilized neural networks to predict dynamics, achieving promising results in simple particle systems. Subsequent studies (sanchez2018graph; kipf2018neural; ajay2018augmenting; li2019propagation; li2019learning; li2019particle) adopted network architectures to the specific features of the systems, thereby enhancing performance. Differentiable simulators are also proposed (degrave2019differentiable; de2018end; hu2019difftaichi), enabling end-to-end training of the physics-based behavior from ground-truth data. The advent of Physics Informed Neural Networks (PINNs) (raissi2019physics; pakravan2021solving) marks a leap forward. These networks incorporate extensive physical information to constrain and guide the learning process, ensuring that predictions adhere more closely to physical laws and has succeeded in domains such as cloths (Geng2020physical) and fluids (um2020solver; gibou2019sharp; chu2022fpinn). Some work (yang2020learning) shifts away from end-to-end structures and use neural networks to optimize part of the simulation. Another line of research generates dynamics through physics-based simulators, where network learns static information while physical laws govern the generation of dynamics (li2023pac; feng2023pie; xie2023physgaussian; feng2024gaussian; jiang2024vr), giving physical meanings to Neural Radiance Fields (NeRF) (mildenhall2020nerf; kerbl20233Dgaussians). These methods demonstrate the benefits of embedding human knowledge into networks to reduce the learning burden.
Background
@@
To make the paper self-contained, we start with a brief review of the preliminaries of elastodynamics and the diffusion model.
Variational optimization of elastodynamics
@@
The dynamic equilibrium of a 3D model can be characterized by $\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L}{\partial\dot{\mathbf{q}}}\right)-\frac{\partial L}{\partial\mathbf{q}}=\mathbf{f}_{q}$, where L=TUL=T-U is system Lagrangian i.e., the difference between the kinematic energy ($T$) and the potential energy ($U$). $\mathbf{q}$ and $\dot{\mathbf{q}}$ are generalized coordinate and velocity. $\mathbf{f}_{q}$ is the generalized external force. With the implicit Euler time integration scheme: $\mathbf{q}^{n+1}=\mathbf{q}^{n}+h\dot{\mathbf{q}}^{n+1}$, $\dot{\mathbf{q}}^{n+1}=\dot{\mathbf{q}}^{n}+h\ddot{\mathbf{q}}^{n+1}$, it can be reformulated as a nonlinear optimization to be solved at each time step:
$\mathbf{q}^{n+1}=\underset{\mathbf{q}}{\mathrm{argmin}}\left\{\frac{1}{2h^{2}}\|\mathbf{q}-\mathbf{\hat{q}}\|^{2}_{\mathbf{M}}+U(\mathbf{q})\right\},$ (1)
@@
where $\mathbf{\hat{q}}=\mathbf{q}^{n}-h\dot{\mathbf{q}}^{n}-h^{2}\mathbf{M}^{-1}\mathbf{f}_{q}$, the superscript $n$ and $n+1$ indicates the time step, $h$ is the time step size, and $\mathbf{M}$ is the mass matrix.
Diffusion model
@@
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{(1-\overline{\alpha}_{t})}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t))+\sigma_{t}\mathbf{z},\mathbf{z}\sim N(\mathbf{0},\mathbf{I}),$ (3)
@@
Methodology
ElastoGen, as illustrated in Fig.1, is a lightweight generative deep model that produces physically grounded 4D content from general object descriptions, such as stiffness or mass. ElastoGen rasterizes the input shape and uses a nested two-level recurrent neural network (RNN) to predict its trajectory sequentially. Each prediction undergoes an accuracy check to ensure physical validity. The network design follows a numerical procedure based on variational optimization as in Eq.(1), ensuring that ElastoGen avoids redundant components that could lead to overfitting. The following sections detail each major module of the pipeline.
@@
Method overview: piece-wise local quadratic approximation
@@
Our elastodynamics generation mimics numerical optimization procedures that minimize the variational energy as in Eq. (1). It is possible to tackle this problem at the global level, i.e., optimizing all the degrees of freedom (DoFs) of the system at once e.g., using Newton’s method. Such a brute-force scheme requires to learn dense inter-correlations among features at all DoFs, which inevitably leads to complex and large-scale network architectures with numerous parameters to be learned.
@@
Alternatively, we opt for a divide-and-conquer way to approach Eq. (1). We consider the total potential energy $U$ as the summation of multiple energies of quadratic form: $U(\mathbf{q})\approx\sum_{i}E_{i}(\mathbf{q}_{i})$, where $E_{i}(\mathbf{q}_{i})=\min_{\mathbf{p}_{i}\in\mathcal{M}_{i}}\frac{\omega_{i}}{2}\left\|\mathcal{G}_{i}[\mathbf{q}_{i}]-\mathbf{p}_{i}\right\|^{2}$. Here, $i$ indicates the $i$-th sub-volume of the object. For instance, one may discretize the object into a tetrahedral mesh, and $E_{i}$ then represents the elastic potential stored at the $i$-th element. $\mathcal{G}_{i}$ denotes a discrete differential operator, which converts positional features $\mathbf{q}_{i}$ to strain-level features. To this end, we build $\mathcal{G}_{i}$ such that $\mathcal{G}_{i}[\mathbf{q}_{i}]=\text{vec}(\mathbf{F}_{i})$, i.e., the vectorized deformation gradient ($\mathbf{F}\in\mathbb{R}^{3\times 3}$) of the sub-volume, which gives the local first-order approximation of the displacement field. The constraint manifold $\mathcal{M}_{i}$ denotes the zero level set of $E_{i}$. In other words, we consider $E_{i}$ as a quadratic energy based on how far local displacement $\mathbf{q}_{i}$ is from its closest energy-free configuration ($\mathbf{p}_{i}$), given the local material stiffness $\omega_{i}$.
@@
Provided the current deformed shape $\mathbf{q}_{i}$, we can find $\arg\min_{\mathbf{p}_{i}}\frac{\omega_{i}}{2}\left\|\mathcal{G}_{i}[\mathbf{q}_{i}]-\mathbf{p}_{i}\right\|^{2}$, which suggests a locally optimal descent direction to reduce $U$. The global displacement can then be obtained by minimizing $\mathbf{q}$ over $E_{i}$ at all the sub-volumes. While this is a global operation, it acts as a Laplacian-like smoothing operator, which can be approximated through repeated local smoothing. This procedure resembles shape matching methods (muller2005meshless) and PD (Bouaziz2014PD), offering a piecewise sequential quadratic programming (SQP) approach (boggs1995sequential) to approximate $U$ locally. ElastoGen functions as a neural version of the aforementioned procedure, with a nested RNN structure that handles local strain relaxation through volume convolutions, ensuring the network remains compact and lightweight.
@@
Unfortunately, real-world materials are more than a collection of quadratic forms. The elastic energy of nonlinear materials cannot be well approximated by a quadratic form of the deformation gradient. This limitation means that shape matching or PD can only handle simplified material behavior. To this end, we augment ElastoGen with a NeuralMTL module to ensure that each local SQP closely matches actual materials.
NeuralMTL & neural projection
@@
The goal of NeuralMTL is to correct local quadratic approximations of $U$ so that ElastoGen faithfully generates physically accurate results for any real-world hyperelastic material. Specifically with NeuralMTL ($\mathcal{N}$), $E_{i}$ becomes:
@@
We set $\omega_{i}$ as $\omega_{i}=V_{i}e$ based on real-world material parameters: Young’s modulus $e$ and the size of the sub-volume $V_{i}$. $\mathcal{G}_{i}$ extracts the deformation gradient $\text{vec}(\mathbf{F}_{i})$ and feeds it to NeuralMTL, $\mathcal{N}$. As the name suggests, $\mathcal{N}$ predicts a neural strain based on the information of local deformation $\mathbf{F}_{i}$. Given the material model and parameters, $\mathcal{N}$ is used for all $E_{i}$, and we do not put a subscript on $\mathcal{N}$. $\|\cdot\|_{F}$ denotes the Frobenius norm. $\mathcal{N}$ predicts a material-space strain prediction, which is then converted to world space by $\mathbf{F}_{i}$. $\mathbf{P}_{i}\in\mathbb{R}^{3\times 3}$ is a rotation matrix i.e., $\mathbf{P}_{i}\in\mathcal{SO}(3)$. Intuitively, as shown in Fig. 2 (a), NeuralMTL warps $\mathbf{F}_{i}$ to a different configuration of $\mathbf{F}_{i}\cdot\mathcal{N}(\mathbf{F}_{i})$ so that the new distance to $\mathbf{P}_{i}$ correctly reflects the local energy landscape of $E_{i}$ as visualized in the inset.
For isotropic elastic materials, we add a nonlinear singular value decomposition (SVD) activation to the operator $\mathcal{G}_{i}$ such that $\mathbf{F}_{i}=\mathbf{U}_{i}\mathbf{S}_{i}\mathbf{V}_{i}^{\top}$. $\mathbf{S}_{i}$ is a diagonal matrix with singular values arranged in descending order, which correspond to the local principal strains. This activation converts $E_{i}$ to:
$\begin{split}E_{i}(\mathbf{q}_{i})=&\frac{\omega_{i}}{2}\|\mathbf{U}_{i}\mathbf{S}_{i}\mathbf{V}_{i}^{\top}\cdot\mathcal{N}(\mathcal{G}_{i}[\mathbf{q}_{i}])-\mathbf{U}_{i}\mathbf{V}_{i}^{\top}\|^{2}_{F}\\ =&\frac{\omega_{i}}{2}\mathrm{tr}\left(\mathbf{S}_{i}\mathbf{S}_{i}\mathbf{V}_{i}^{\top}\cdot\mathcal{N}(\mathcal{G}_{i}[\mathbf{q}_{i}])\cdot\mathcal{N}^{\top}(\mathcal{G}_{i}[\mathbf{q}_{i}])\mathbf{V}_{i}\right.\\ &\left.+\mathbf{I}-2\mathbf{V}_{i}\mathbf{S}_{i}\mathbf{V}_{i}^{\top}\cdot\mathcal{N}(\mathcal{G}_{i}[\mathbf{q}_{i}])\right).\end{split}$ (5)
@@
$\begin{split}&E_{i}=\frac{\omega_{i}}{2}\mathrm{tr}\left(\mathbf{Q}_{i}\mathbf{Q}_{i}^{\top}\right)+\frac{3\omega_{i}}{2}-\omega_{i}\ \mathrm{tr}\left(\mathbf{Q}_{i}\right),\\ &\mathbf{Q}_{i}(\mathbf{S}_{i})=\mathbf{S}_{i}\left(\mathbf{N}_{i}+\mathbf{N}_{i}^{\top}\right).\end{split}$ (7)
Intuitively, this activation escalates the order of the neural strain predicted by $\mathcal{N}$, pushing it to become a nonlinear strain estimation with a prescribed format — just like upgrading an infinitesimal strain to Green’s strain to better measure large rotational deformation. As a result, the neural projection corresponding to our NeuralMTL can be easily obtained as $\mathbf{P}_{i}=\mathbf{U}_{i}\mathbf{V}_{i}^{\top}$, i.e., the rotational component from $\mathbf{F}_{i}$. This is an important property of NeuralMTL — if we choose to employ the network to learn an adjustment of $\mathbf{P}_{i}$ (which is also technically feasible), the local relaxation that predicts $\mathbf{P}_{i}$ becomes complicated, and the generation is less robust.
Given an input 3D object, ElastoGen rasterizes it into a set of sub-volumes. For a user-specified sub-volume, which in our implementation is a voxel that intersects with the object, $\mathcal{G}_{i}$ operator extracts the local covariance matrix of the displacement field over this sub-volume. Let $\mathbf{A}_{i}=[\mathbf{q}_{1},\mathbf{q}_{2},...\mathbf{q}_{k}]\in\mathbb{R}^{3\times k}$ and $\bar{\mathbf{A}}_{i}=[\bar{\mathbf{q}}_{1},\bar{\mathbf{q}}_{2},...\bar{\mathbf{q}}_{k}]$ be deformed and rest-shape position of vertices of a sub-volume with $k$ vertices ($k=8$ for a cubic volume). $\mathcal{G}_{i}$ has an analytic format of: